OpenStudy (anonymous):

A person standing 151 centimeters from a mirror notices that the angle of depression from his eyes to the bottom of the mirror is 11°, while the angle of elevation to the top of the mirror is 11°. Find the vertical dimension of the mirror. (Round your answer to the nearest whole number.)

OpenStudy (anonymous):

tan(11)=(h/2)/151, therefore h=2*tan(11)*151

OpenStudy (anonymous):

Would you please be able to show me what the picture would look like? That'd make it easier for me to understand your work.